OpenStudy (anonymous):

Find the simple interest and the total amount due on the loan. $600 is the Principal, 3% is the rate, 3 months for time.

OpenStudy (zale101):

A = P(1 + rt) <<----Simple Interest Equation Where: A is the Total Amount due P is the Principal Amount <-- $600 I is the Interest Amount <--usually 1 r is the Rate of Interest per year in decimal; so r = R/100 <<--3$/100=0.03 t is the Time Period involved in months or years, you have 3 months, so that means \(\frac{ 3~months }{ 12~months/year }=0.25~years\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The simple interest I got is 4.5

OpenStudy (anonymous):

@zale101 is the simple interest right

OpenStudy (zale101):

Can you please show your work. That is not the correct answer.

OpenStudy (anonymous):

4.50

OpenStudy (anonymous):

is 4.50 the simple interest